OpenStudy (anonymous):

One leg of a right triangle measures 8 inches and the hypotenuse 17 inches. Determine the length of the other leg. a. 9 inches b. 13 inches c. 15 inches d. 17 inches

OpenStudy (anonymous):

The Pythagorean Theorem says that for right triangles, a^2 + b^2 = c^2. 8^2 = 64, and 17^2 = 289. Subtract 289 - 64. What do you get? @hellokitty32695

OpenStudy (anonymous):

You get 225, and the square root of 225 = 15.

OpenStudy (anonymous):

Don't give answers. Read the CoC. Right @iGreen ?